Kumi is 4 times as old as Aimee. 3 years ago Kumi was 5 times as old as Aimee. How old is Aimee today?
A)9 B)12 C)15 D)45 E)48

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52788)

.
Kumi is 4 times as old as Aimee. 3 years ago Kumi was 5 times as old as Aimee. How old is Aimee today?
A)9 B)12 C)15 D)45 E)48
~~~~~~~~~~~~~~~~

 x = Aimee's age, in years.

4x = Kumi's age


Equation

    4x - 3 = 5(x-3)


Solution

    4x - 3 = 5x - 15

    15 - 3 = 5x - 4x

       12  =    x


ANSWER.  Aimee is 12 years old.

Solved.

Check it on your own that my answer is correct.



Answer by greenestamps(13200)



If a formal algebraic solution is required, then something like the other tutor provided is a good example.

If the speed of getting an answer is important and formal algebra is not required, note that, because 5 distinct answer choices are given, the problem can be solved most quickly simply by trying the answer choices.

Doing that can also give you some good mental exercise.

Kumi is 4 times as old as Aimee, and the answer choices are for Aimee's age. So note immediately that answer choices D and E are unreasonable, since they would make Kumi 180 or 192 years old.

Try answer choice A:
Aimee is 9, so Kumi is 36
3 years ago their ages were 6 and 33
No, 33 is not 5 times 6

Try answer choice B:
Aimee is 12, so Kumi is 4*12=48
3 years ago their ages were 9 and 45
Yes, 45 is 5 times 9

ANSWER: B)12